OpenStudy (anonymous):

1. Write an equation to solve the following. Show all your work, including the steps for solving the equation and checking your answer. -4 times a number plus 16 is -116. Find the number

OpenStudy (anonymous):

36

OpenStudy (anonymous):

-4 * x + 16 = -116 -4x+ 16 = -116 then subtract -16 on both sides of the equation. the 16 cancels and -116 - -16= -132 Your next line should have -4x = 132 then you divide -4 on both sides which would leave you with x= 36

OpenStudy (anonymous):